# Task-conditioned adaptation of visual features in multi-task policy learning

### v1

@@ Abstract
Successfully addressing a wide variety of tasks is a core ability of autonomous agents, which requires flexibly adapting the underlying decision-making strategies and, as we argue in this work, also adapting the underlying perception modules. An analogical argument would be the human visual system, which uses top-down signals to focus attention determined by the current task. Similarly, in this work, we adapt pre-trained large vision models conditioned on specific downstream tasks in the context of multi-task policy learning. We introduce task-conditioned adapters that do not require finetuning any pre-trained weights, combined with a single policy trained with behavior cloning and capable of addressing multiple tasks. We condition the policy and visual adapters on task embeddings, which can be selected at inference if the task is known, or alternatively inferred from a set of example demonstrations. To this end, we propose a new optimization-based estimator. We evaluate the method on a wide variety of tasks of the CortexBench benchmark and show that, compared to existing work, it can be addressed with a single policy. In particular, we demonstrate that adapting visual features is a key design choice and that the method generalizes to unseen tasks given visual demonstrations.
Figure 1: Task-conditioned adaptation: We show that a single policy can be trained to address multiple heterogenous tasks including manipulation, legged motion etc., and that few-shot learning is possible to address tasks given as demonstration but unseen during training. A key element is the task-conditioned adaptation of visual features.
@@ 1 Introduction
We propose task-conditioned adaptation, which allows leveraging the high-quality representations of generally pre-trained large vision models, while keeping the required flexibility to address a wide variety of tasks, and also new (unseen) tasks. We introduce a set of task-conditioned visual adapters that can be inserted inside a pre-trained visual Transformer [33]-based backbone. The task is characterized by an embedding space, which is learned from supervision during training. We show that this embedding space captures regularities of tasks and demonstrate this with few-shot capabilities: the single policy and (adapted) visual representations can address new unseen tasks, whose embedding is estimated from a few demonstrations (cf. Fig. 1).
@@
Refer to caption (b) Optimization of the task embedding for the Few-shot  setting
@@
Figure 3: Method overview: (a) the adapted policy is trained with behavior cloning from expert demonstrations and given a visual encoder pre-trained with masked auto-encoding. The model is conditioned on a task-embedding learned from GT 1-in-K task identifiers. (b) In the Few-shot  case, a task embedding is estimated by optimization, maximizing the likelihood of a given demonstration of an unknown task. (c) Inference uses a task embedding given in the Known task  case, or estimated in the Few-shot  case.
@@ 2 Related Work
Pre-trained visual representations for robotics — Backbone models pre-trained on large and diverse data have shown great promises in NLP [14, 2, 24, 18], CV [26, 22, 9, 5], and more recently, robotics [20, 23, 21, 27, 19, 16, 35, 36]. Parisi et al. [23] study visual pre-training methods for visuo-motor control, showing the quality of self-supervised representations. Nair et al. [21] introduce R3M, a general vision model pre-trained on egocentric video data to capture temporal dynamics and semantic features, improving downstream manipulation performance. Radosavovic et al. [27] employ the well-known MAE framework [9] to pre-train a single vision encoder applied to robots in the real world. Ma et al. [19] pre-train a visual model with a self-supervised value function objective on egocentric human videos to improve the training of control policies. Finally, recent work [16, 35, 36] has shown the great promise of pre-trained models, either CLIP-based [16] or self-supervised [35, 36], in visual navigation.
@@
While prior work learns a specific set of adapters for each task, we argue that tasks share similarities and explore these regularities through a single multi-task policy.
@@ 3.1 Base agent architecture
Visual encoder without adapters — following [20], the visual encoder, denoted as $\phi$, is a ViT model [7], which has been pre-trained with masked auto-encoding (MAE). We keep pre-trained weights from VC-1 in [20], which are publicly available. However, we change the way the representation is collected from the pre-trained model. Unlike [20], the representation is not taken as the embedding of the ’CLS’ token, which we consider to be undertrained by the MAE pretext task. Instead, we train a fully-connected layer $\psi$ to aggregate all the token representations of the last layer of the ViT except the ’CLS’ token. The visual observation $\mathbf{v}_{t}$ associated with timestep $t$ is thus encoded as
@@
where ${\theta}_{\phi}$ and ${\theta}_{\psi}$ are weights parametrizing $\phi$ and $\psi$, respectively. $\mathbf{r}_{t}\in\mathbb{R}^{3\times d_{r}}$ as it contains the $d_{r}$-dim encoding of each of the $3$ last visual frames processed as a data batch, where $d_{r}$ is the output dimension of $\psi$.
As this will become relevant later, we recall here that a ViT $\phi$ is composed of a sequence of $N_{s}$ self-attention blocks, where $\phi_{i}$ is the layer at index $i$. If we denote $\mathbf{s}^{i}_{t}$ as the internal hidden representation predicted at layer $i$, we have,
$\mathbf{s}^{i}_{t}=\phi_{i}(\mathbf{s}^{i-1}_{t}))$ (2)
@@
Table 1: Known task  — Impact of visual adapters: Validation and test performance on known tasks of different baselines highlighting the gain brought by adapters. Both visual and top adapters bring a boost in performance, and conditioning them on the learned task embedding increases performance. Our proposed architecture involving a multi-task policy outperforms single-task policies with VC-1 non-adapted features. MT $\pi$: multi-task policy – NC: Non-conditioned – C: Conditioned – Task emb.: whether to input at evaluation time, either the learned task embedding and chosen from ground-truth (L), a random vector as the task embedding (Rd), or a randomly picked task embedding among the set of $K{=}12$ embeddings (RdP) – Benchmarks avg: average performance accross the $3$ considered benchmarks (Adroit, DMC, MetaWorld) – Tasks avg: average performance across all $12$ known tasks.
@@
Single-task policy — again following [20], the policy $\pi$ is memory-less and implemented as an MLP predicting actions from the input which is a concatenation of the current frame, two frame differences and the proprioception input $\mathbf{p_{t}}$,
@@ 3.2 Adaptation
Our key contributions are visual adapter modules along with a multi-task policy, which are all conditioned on the task at hand. This is done with a specific task embedding for each task, taken from an embedding space of dimension $d_{e}$, which is aimed to have sufficient regularities to enable few-show generalization to unseen tasks. Importantly, the different adapters and the multi-task policy are conditioned on the same task embedding, leading to a common and shared embedding space. For the known task setting, where the ground-truth label of the task is available, the task embedding is projected from a 1-in-K vector with a linear function $g$ trained jointly with the adapters and the policy with the downstream loss (imitation learning). In the Few-shot setting, at test time a new unknown task is described with a few demonstrations, and a task embedding is estimated through optimization, as will be detailed in subsection 3.4. Figure 3 outlines the architecture, its details will be given in the supplementary material.
@@
Middle adapters — we add one trainable adapter after each such ViT block to modulate its output. We introduce a set of middle adapters $A=\{\alpha_{i}\}_{[i=1..N_{s}]}$, where $\alpha_{i}$ is a 2-layer MLP. In the modified visual encoder $\phi^{m}$, each adapter modulates the output of the corresponding self-attention block and is conditioned on the task embedding $\mathbf{e}$. The output of a middle adapter is combined with the one of the self-attention layer through a residual connection. If we denote $\mathbf{s}^{m,i}_{t}$ as the internal hidden representation predicted at layer $i$, the associated forward pass of a given layer becomes,
$\mathbf{s}^{m,i}_{t}$ $=\phi^{m}_{i}(\mathbf{s}^{m,(i-1)}_{t})$ (4)
$=\phi_{i}(\mathbf{s}^{m,(i-1)}_{t})+\alpha_{i}(\phi_{i}(\mathbf{s}^{m,(i-1)}_{t}),\mathbf{e})$ (5)
$=\mathbf{s}^{i}_{t}+\alpha_{i}(\mathbf{s}^{i}_{t},\mathbf{e})$ (6)
@@
$\mathbf{r}_{t}^{m}=\tau\Bigr{[}\psi\bigr{[}\phi^{m}(\mathbf{v}_{t},\mathbf{e};{\theta}_{\phi},{\theta}_{A});{\theta}_{\psi}\bigr{]},\mathbf{e};\theta_{\tau}\Bigr{]}$ (7)
@@
Figure 5: Known task  — Per-task performance: performance of single-task policies (row (a) in Table 1), our approach without any adapter (row (b) in Table 1) and with conditioned middle and top adapters (row (f) in Table 1). The proposed adapters lead to a performance gain on most tasks, and our multi-task solution is competitive with single-task policies.
@@ 3.3 Training
We train the model by keeping the weights of the pre-trained vision-encoder model $\theta_{\phi}$ frozen, only the weights of the adapter modules (${\theta}_{A}$, ${\theta}_{\tau}$), aggregation layer (${\theta}_{\psi}$), embedding layer ($\theta_{g}$) and multi-task policy (${\theta}_{\pi^{m}}$) are trained, cf. Fig. 3a. Lets’ denote by $\Theta{=}{\{\theta}_{A},{\theta}_{\tau},{\theta}_{\psi},{\theta}_{g},{\theta}_{\pi^{m}}\}$ the set of optimized weights. We train with imitation learning, more specifically Behavior Cloning (BC): for each known task $t_{i}^{k}$, we have access to a set of $N_{i}$ expert trajectories that are composed of $T_{i}$ discrete steps, including expert actions. The optimization problem is given as
$\hat{\Theta}=\arg\min_{\Theta}\sum_{i=1}^{K}\sum_{n=1}^{N_{i}}\sum_{t=1}^{T_{i}}\mathcal{L}(\hat{\mathbf{a}}_{t},{\mathbf{a}}_{t}^{i,n})$ (9)
where ${\mathbf{a}}_{t}^{i,n}$ is the ground-truth expert action for a given step in a trajectory, and $\mathcal{L}$ is the Mean Squared Error loss.
@@ 3.4 Few-shot adaption to new tasks
For the Few-shot  setting, the task embedding $\mathbf{e}$ is unknown at inference and needs to be estimated from a set of $N_{d}$ example demonstrations $\mathcal{D}=\{d_{i}\}_{i=1..N_{d}}$ where $d_{i}=\{(\mathbf{v}^{i*}_{t},\mathbf{p}^{i*}_{t},\mathbf{a}^{i*}_{t})\}_{t=1..T}$ is composed of observations and actions. We exploit the conditioning property of the policy itself to estimate the embedding $\hat{\mathbf{e}}$ as the one which obtains the highest probability of the demonstration actions, when the policy is applied to the demonstration inputs, i.e.
$\hat{\mathbf{e}}=\arg\min_{\mathbf{e}}\sum_{i}\sum_{t}\mathcal{L}(\pi^{m}\left(\left[\mathbf{\tilde{r}}^{mi*}_{t},\mathbf{e}\right],\theta_{\pi^{m}},\right),{\mathbf{a}}_{t}^{i*}),$ (10)
where $\mathbf{\tilde{r}}^{mi*}$ is the representation extracted from the demonstration input $(\mathbf{v}^{i*}_{t},\mathbf{p}^{i*}_{t})$, and which itself depends on $\mathbf{e}$ (this is not made explicit in the notation). The minimization is carried out with SGD from an embedding initialized to zero.
Figure 6: Few-shot  — Per-task performance on unknown tasks: After optimizing the task embedding independently for each task from 5 demonstrations, our method can adapt to many of them, without requiring any weight finetuning.
@@ 4 Experiments
Evaluation — to better handle possible overfit on hyper-parameter selection, our evaluation setup is slightly different from [20] as we perform $100$ validation rollouts to select the best checkpoint of each model, and then test the chosen model on $100$ test rollouts. For our multi-task policy, the best checkpoint is the one with the highest average validation performance across all tasks. Single-task policies are validated only on the task they were trained on, giving them an advantage, and for this reason, they are reported as “soft upper bounds”. We report the average performance and standard deviation among $3$ trained models (3 random seeds) for each variant.
@@
Evaluation metrics — Following [20], we consider a rollout success ($1$ if the task was completed properly, 00 otherwise) for tasks in the Adroit and MetaWorld benchmarks, and report the agent’s normalized return for Deepmind Control. Performance is averaged across rollouts. Episodes have a maximum length of $1000$ steps and each step reward is comprised between 00 and $100$ in DMC, normalization is therefore done by dividing the agent’s return by $10$.
@@
— Impact of visual adapters — Table 1 presents a detailed comparison of different methods on the known task setting. The baseline in row (a) follows the setup in [20] to train single-task policies (one for each task) from non-conditioned VC-1 features. For this variant, we also use the representation of the ’CLS’ token as the vector fed to the policy as done to respect its original choice, while all other baselines use our proposed token aggregation layer.
Row (b) is our multi-task policy without any adapter, as expected there is some performance drop when compared to the specialized policies in row (a), as the problem to solve has become more difficult. Adding adapters and conditioning them on the task embedding, shown in rows (c)-(f), brings a boost in performance, both for middle and for top adapters. In particular, conditioning adds a further boost compared to non-conditioned adapters, with all choices enabled, row (f), obtaining the best average performance.
Rows (g) and (h) are ablation experiments that evaluate the impact of choosing completely random task embeddings, row (g), or of taking a random choice between the 12 learned embeddings, row (h). In both cases, the performance collapses.
@@
Figure 7: Few-shot  — Qualitative results on unknown tasks: Three successful policy rollouts on unknown tasks. The multi-task approach tackles these new tasks involving different objects or manipulation requirements than what it was trained on.
Figure 8: Few-shot  — Failure cases on unknown tasks: The first row (button-press-wall task) shows an example where the method performs the task correctly until the end where it fails to properly push the button fully. This behavior seems to happen often for this button-press-wall task, explaining its $0\%$ success despite being able to reproduce most of the behavior. In the second row (push-back task), the policy is able to properly move the cube but fails to bring it to the goal position (green dot).
@@
— Additional ablation studies — are presented in the appendix (Sections B, C). Section B shows that conditioning the policy when using task-conditioned adapters is not necessary, that our aggregation layer works better than using the ’CLS’ token as input to the policy and further highlights the impact of both middle and top adapters. Section C shows that our adapters improve visual embeddings extracted by two other state-of-the-art pre-trained backbones, i.e. PVR [23] and MVP [27], confirming the conclusions drawn from experiments on VC-1.
@@
— adaptation to new tasks without finetuning model weights is an important ability of any general policy, which we evaluate with the following experiments: for each task within a set of unknown tasks (cf. Figure 2), we collect only $5$ demonstrations used to optimize a task embedding specific to this task with the method detailed in section §3.4. We then test the method conditioned on the optimized embedding on $100$ test rollouts.
@@
To generate the set $T^{u}$ of unknown tasks, we select tasks from the MetaWorld dataset which had not been selected for the CortexBench benchmark, and which are therefore not part of the set of known tasks $T^{k}$ on which we trained the policy and adapters. Unfortunately for these tasks, there do not exist expert demonstrations as they have been defined as environments for RL training only. We collect demonstrations using single-task policies from TD-MPC2 [8] that were specifically trained on each task of MetaWorld independently. To ensure high-quality demonstrations, we only consider tasks where TD-MPC2 policies reach a success rate higher than $95\%$. Furthermore, to be compatible with the choices of the CortexBench authors, in particular, to keep the same camera locations, we filter out some of the tasks where the goal is not always visible in the camera FOV. This leads to a set of $15$ unknown tasks that are quite different from the tasks in the training set $T^{k}$ as they involve different objects (handle press, faucet, plate, door, window, etc.) and types of manipulation (sliding an object, lowering a press, opening a window, etc.). Each collected demonstration is a sequence of visual frames, proprioception inputs, and expert actions. The optimization of the task embedding is performed independently for each task (cf. §3.4). We use the AdamW optimizer and a learning rate of $1e{-1}$ during the task embedding search.
@@
Figure 6 presents the per-task performance in this setting. Despite the large variations in terms of objects and types of manipulations between the new tasks in $T^{u}$ and the ones in the training set ($T^{k}$), the proposed multi-task policy can adapt to many of them, without requiring any weight finetuning. Interestingly, the method performs particularly well on the Drawer Close task, as the time-inversed Drawer Open task was in the training set. This provides some evidence that the method can exploit regularities between tasks, which seem to be captured by the task embedding space, making it possible to generalize to unseen variations. Figure 7 shows qualitative examples of successful rollouts on the new tasks. The multi-task policy is able to manipulate new objects (faucet, plate, window) that were not encountered at training time. It also successfully performs diverse new moves such as rotating the faucet or sliding the plate.
@@
Finally, Figure 8 shows failure cases of our method on unknown tasks. As can be seen on the first row, the policy avoids the wall, reaches the button, and starts pushing it, but fails to push it fully. This particular behavior was also observed on other rollouts, explaining the $0\%$ success on this button-push-wall task while mastering a part of the required sub-skills. On the second row, the policy is being evaluated on the push-back task. It is able to move the cube but fails to bring it to the goal location (green dot). This gives some indication of the difficulty of the few-shot generalization case: exploiting regularities in the task space requires that tasks be performed more than just approximately, as often the success metric is sparse, and rollouts only count to the metric when they are executed fully and correctly.

### v2

@@ Abstract
Successfully addressing a wide variety of tasks is a core ability of autonomous agents, requiring flexibly adapting the underlying decision-making strategies and, as we argue in this work, also adapting the perception modules. An analogical argument would be the human visual system, which uses top-down signals to focus attention determined by the current task. Similarly, we adapt pre-trained large vision models conditioned on specific downstream tasks in the context of multi-task policy learning. We introduce task-conditioned adapters that do not require finetuning any pre-trained weights, combined with a single policy trained with behavior cloning and capable of addressing multiple tasks. We condition the visual adapters on task embeddings, which can be selected at inference if the task is known, or alternatively inferred from a set of example demonstrations. To this end, we propose a new optimization-based estimator. We evaluate the method on a wide variety of tasks from the CortexBench benchmark and show that, compared to existing work, it can be addressed with a single policy. In particular, we demonstrate that adapting visual features is a key design choice and that the method generalizes to unseen tasks given a few demonstrations.
Figure 1: Task-conditioned adaptation: A single policy can be trained to address multiple heterogenous tasks including manipulation, legged motion etc., and few-shot learning is possible to address tasks given as demonstrations but unseen during training. A key element is the task-conditioned adaptation of visual features.
@@ 1 Introduction
We propose task-conditioned adaptation, which allows leveraging the high-quality representations of generally pre-trained large vision models, while keeping the required flexibility to address a wide variety of tasks, and also new (unseen) tasks. We introduce a set of task-conditioned visual adapters that can be inserted inside a pre-trained visual Transformer [33]-based backbone. The task is characterized by an embedding space, which is learned from supervision during training. We show that this embedding space captures regularities of tasks and demonstrate this with few-shot capabilities: the single policy and (adapted) visual representations can address new unseen tasks, whose embedding is estimated from a few demonstrations (cf. Figure 1).
@@
Refer to caption (b) Optimization of the task embedding in the Few-shot  setting.
@@
Figure 3: Method overview: (a) the adapted policy is trained with behavior cloning from expert demonstrations and given a visual encoder pre-trained with MAE. The model is conditioned on a task embedding learned from ground-truth 1-in-K task identifiers. (b) In the Few-shot  case, a task embedding is estimated by optimization, maximizing the likelihood of a given demonstration of an unknown task. (c) Inference uses a task embedding given in the Known task  case, or estimated in the Few-shot  case.
@@ 2 Related Work
Pre-trained visual representations for robotics — Backbone models pre-trained on large and diverse data have shown great promises in NLP [14, 2, 24, 18], CV [26, 22, 9, 5], and more recently, robotics [20, 23, 21, 27, 19, 16, 35, 36]. Parisi et al. [23] study visual pre-training methods for visuo-motor control, showing the quality of self-supervised representations. Nair et al. [21] introduce R3M, a general vision model pre-trained on egocentric video data to capture temporal dynamics and semantic features, improving downstream manipulation performance. Radosavovic et al. [27] employ the well-known MAE framework [9] to pre-train a single vision encoder applied to robots in the real world. Ma et al. [19] pre-train a visual model with a self-supervised value function objective on egocentric human videos to improve control policies. Finally, recent work [16, 35, 36] has shown the great promise of pre-trained models, either CLIP-based [16] or self-supervised [35, 36], in visual navigation.
@@
While prior work learns a specific set of adapters for each task, we argue that tasks share similarities and explore these regularities with a single set of task-conditioned adapters.
@@ 3.1 Base agent architecture
Visual encoder without adapters — following [20], the visual encoder, denoted as $\phi$, is a ViT model [7] pre-trained with masked auto-encoding (MAE). We keep pre-trained weights from VC-1 in [20], which are publicly available. However, we change the way the representation is collected from the pre-trained model. Unlike [20], the representation is not taken as the embedding of the ’CLS’ token, which we consider to be undertrained by the MAE pretext task. Instead, we train a fully-connected layer $\psi$ to aggregate all the token representations of the last layer of the ViT except the ’CLS’ token. The visual observation $\mathbf{v}_{t}$ associated with timestep $t$ is thus encoded as
@@
where ${\theta}_{\phi}$ and ${\theta}_{\psi}$ are weights parametrizing $\phi$ and $\psi$ respectively. $\mathbf{r}_{t}\in\mathbb{R}^{3\times d_{r}}$ as it contains the $d_{r}$-dim encoding of each of the $3$ last visual frames processed as a data batch, where $d_{r}$ is the output dimension of $\psi$.
As this will be relevant later, we recall here that a ViT $\phi$ is composed of a sequence of $N_{l}$ self-attention blocks, where $\phi_{l}$ is the layer at index $l$. If we denote the internal hidden representation predicted at layer $l$ as $\mathbf{s}^{l}_{t}$, and omit the weights of $\phi_{l}$ for simplicity, we have,
$\mathbf{s}^{l}_{t}=\phi_{l}(\mathbf{s}^{l-1}_{t}),$ (2)
@@
Table 1: Known task  — Impact of visual adapters: Validation and test performance on known tasks of different baselines highlighting the gain brought by adapters. Both middle and top adapters bring a boost in performance, and conditioning them on the learned task embedding increases performance. Our multi-task policy outperforms single-task policies with VC-1 non-adapted features. MT $\pi$: multi-task policy – NC: Non-conditioned – C: Conditioned – Task emb.: whether to input at evaluation time, either the learned task embedding and chosen from ground-truth (L), a random vector as the task embedding (Rd), or a randomly picked task embedding among the set of $K{=}12$ embeddings (RdP) – Benchmarks avg: average performance across the $3$ considered benchmarks (Adroit, DMC, MetaWorld) – Tasks avg: average performance across all $12$ known tasks. Performance is reported as mean $\pm$ std over $3$ training runs (seeds).
@@
Single-task policy — following [20], the policy $\pi$ is implemented as an MLP predicting actions from the input which is a concatenation of the current frame, two frame differences and the proprioception input $\mathbf{p_{t}}$,
@@ 3.2 Adaptation
Our key contributions are visual adapter modules along with a multi-task policy, which are all conditioned on the task at hand. This is done with a specific task embedding for each task, taken from an embedding space of dimension $d_{e}$, which is aimed to have sufficient regularities to enable few-show generalization to unseen tasks. Importantly, the different adapters and the multi-task policy are conditioned on the same task embedding, leading to a common and shared embedding space. For the known task setting, where the ground-truth label of the task is available, the task embedding is projected from a 1-in-K vector with a linear function $g$ trained jointly with the adapters and the policy with the downstream loss (imitation learning). In the Few-shot setting, at test time a new unknown task is described with a few demonstrations, and a task embedding is estimated through optimization, as will be detailed in subsection 3.4. Figure 3 outlines the architecture, its details will be given in the appendix.
@@
Middle adapters — we add one trainable adapter after each ViT block to modulate its output. We introduce a set of middle adapters $A=\{\alpha_{l}\}_{[l=1..N_{l}]}$, where $\alpha_{l}$ is a 2-layer MLP. In the modified visual encoder $\phi^{m}$, each adapter modulates the output of the corresponding self-attention block and is conditioned on the task embedding $\mathbf{e}$. Its output is combined with the one of the self-attention layer through a residual connection. If we denote the internal hidden representation predicted at layer $l$ as $\mathbf{s}^{m,l}_{t}$, and omit references to the weights of $\phi_{l}$ and $\alpha_{l}$ as in eq. (2) for simplicity, the associated forward pass of a given layer becomes,
$\mathbf{s}^{m,l}_{t}$ $=\phi^{m}_{l}(\mathbf{s}^{m,(l-1)}_{t})$ (4)
$=\phi_{l}(\mathbf{s}^{m,(l-1)}_{t})+\alpha_{l}(\phi_{l}(\mathbf{s}^{m,(l-1)}_{t}),\mathbf{e}).$ (5)
@@
$\mathbf{r}_{t}^{m}=\tau\Bigr{[}\psi\bigr{[}\phi^{m}(\mathbf{v}_{t},\mathbf{e};{\theta}_{\phi},{\theta}_{A});{\theta}_{\psi}\bigr{]},\mathbf{e};\theta_{\tau}\Bigr{]},$ (6)
@@
Figure 5: Known task  — Per-task performance of policies in Table 1: single-task policies (row (a)), our approach without any adapter (row (b)) and with conditioned middle and top adapters (row (f)). The adapters lead to a performance gain on most tasks, and our multi-task solution is competitive with single-task policies. Colored bars and error bars respectively show mean and std over $3$ training runs (seeds).
@@ 3.3 Training
We train the model by keeping the weights of the pre-trained vision-encoder model $\theta_{\phi}$ frozen, only the weights of the adapter modules (${\theta}_{A}$, ${\theta}_{\tau}$), aggregation layer (${\theta}_{\psi}$), embedding layer ($\theta_{g}$) and multi-task policy (${\theta}_{\pi^{m}}$) are trained, cf. Figure 3a. Lets’ denote by $\Theta{=}{\{\theta}_{A},{\theta}_{\tau},{\theta}_{\psi},{\theta}_{g},{\theta}_{\pi^{m}}\}$ the set of optimized weights. We train with imitation learning, more specifically Behavior Cloning (BC): for each known task $t_{i}^{k}$, we have access to a set of $N_{i}$ expert trajectories that are composed of $T_{i}$ discrete steps, including expert actions. The optimization problem is given as
$\hat{\Theta}=\arg\min_{\Theta}\sum_{i=1}^{K}\sum_{n=1}^{N_{i}}\sum_{t=1}^{T_{i}}\mathcal{L}(\hat{\mathbf{a}}_{t}^{i,n},{\mathbf{a}}_{t}^{i,n}),$ (8)
where $\hat{\mathbf{a}}_{t}^{i,n}$ and ${\mathbf{a}}_{t}^{i,n}$ are the predicted and ground-truth actions for a given step in a trajectory, and $\mathcal{L}$ is the Mean Squared Error loss.
@@ 3.4 Few-shot adaption to new tasks
For the Few-shot  setting, the task embedding $\mathbf{e}$ is unknown at inference and needs to be estimated from a set of $N_{d}$ example demonstrations $\mathcal{D}{=}\{d_{n}\}_{[n=1..N_{d}]}$ where $d_{n}{=}\{(\mathbf{v}^{n*}_{t},\mathbf{p}^{n*}_{t},\mathbf{a}^{n*}_{t})\}_{[t=1..T_{d}]}$ is composed of observations and actions, with $T_{d}$ being the length of each demonstration. We exploit the conditioning property of the policy itself to estimate the embedding $\hat{\mathbf{e}}$ as the one which obtains the highest probability of the demonstration actions, when the policy is applied to the demonstration inputs, i.e.
$\hat{\mathbf{e}}=\arg\min_{\mathbf{e}}\sum_{n=1}^{N_{d}}\sum_{t=1}^{T_{d}}\mathcal{L}(\pi^{m}\left(\left[\mathbf{\tilde{r}}^{mn*}_{t},\mathbf{e}\right],\theta_{\pi^{m}}\right),{\mathbf{a}}_{t}^{n*}),$ (9)
where $\mathbf{\tilde{r}}^{mn*}$ is the representation extracted from the demonstration input $(\mathbf{v}^{n*}_{t},\mathbf{p}^{n*}_{t})$, and which itself depends on $\mathbf{e}$ (not made explicit in the notation). The minimization is carried out with SGD from an embedding initialized to zero.
Figure 6: Few-shot  — Per-task performance: After optimizing the task embedding for each task from 5 demonstrations, our method can adapt to many of them (without finetuning). Colored bars and error bars respectively show mean and std over $3$ training runs (seeds).
@@ 4 Experiments
Evaluation — to better handle possible overfit on hyper-parameter selection, our evaluation setup is slightly different from [20] as we perform $100$ validation rollouts to select the best checkpoint of each model, and then test the chosen model on $100$ test rollouts. For our multi-task policy, the best checkpoint is the one with the highest average validation performance across all tasks. Single-task policies are validated only on the task they were trained on, giving them an advantage, and for this reason, they are reported as “soft upper bounds”. We report the average performance and standard deviation among $3$ trained models (3 random seeds) for each variant as mean $\pm$ std (Table 1).
@@
Evaluation metrics — Following [20], we consider a rollout success ($1$ if the task was completed properly, 00 otherwise) for tasks in the Adroit and MetaWorld benchmarks, and report the normalized return for DMC. Episodes have a maximum length of $1000$ steps and each step reward is comprised between 00 and $100$ in DMC, normalization is therefore done by dividing the agent’s return by $10$. For all tasks, performance is averaged across rollouts.
@@
— Impact of visual adapters — Table 1 presents a detailed comparison of different methods on the known task setting. The baseline in row (a) follows the setup in [20] to train single-task policies (one per task) from non-conditioned VC-1 features. For this variant, we use the representation of the ’CLS’ token as the vector fed to the policy as done in [20], while all other baselines use our proposed token aggregation layer.
Row (b) is our multi-task policy without any adapter. As expected there is a performance drop compared to the specialized policies in row (a), as the problem to solve has become more difficult. Adding adapters and conditioning them on the task embedding, shown in rows (c)-(f), brings a boost in performance, both for middle and top adapters. In particular, conditioning adds a further boost compared to non-conditioned adapters, with all choices enabled, row (f), obtaining the best average performance.
Rows (g) and (h) are ablation experiments evaluating the impact of choosing random task embeddings, row (g), or of taking a random choice between the 12 learned embeddings, row (h). In both cases, the performance collapses.
@@
(a) Successful rollouts
(b) Failed rollouts
Figure 7: Few-shot  — Qualitative results: (a) The policy tackles new tasks involving objects and/or manipulation requirements unseen during training. (b) In the first row (button-press-wall task), it performs the task correctly until the end where it fails to properly push the button fully. In the second row (push-back task), it properly moves the cube but fails to bring it to the goal position (green dot).
@@
— Additional ablation studies — are presented in the appendix (Sections B and C). Section B shows that conditioning the policy when using task-conditioned adapters is not necessary, that our aggregation layer works better than using the ’CLS’ token as input to the policy and further highlights the impact of both middle and top adapters. Section C shows that our adapters improve visual embeddings extracted by two state-of-the-art pre-trained backbones, i.e. PVR [23] and MVP [27], confirming the conclusions drawn from experiments on VC-1.
@@
— adaptation to new tasks without finetuning model weights is an important ability of any general policy, which we evaluate with the following experiments: for each task within a set of unknown tasks (cf. Figure 2), we collect only $5$ demonstrations used to optimize a task embedding specific to this task with the method detailed in section §3.4. We then evaluate the method conditioned on the optimized embedding on $100$ test rollouts.
@@
To generate the set $T^{u}$ of unknown tasks, we select tasks from the MetaWorld dataset that do not belong to CortexBench, and are thus not part of the set of training known tasks $T^{k}$. We collect demonstrations using single-task policies from TD-MPC2 [8] that were specifically trained on each task of MetaWorld independently. To ensure high-quality demonstrations, we only consider tasks where TD-MPC2 policies reach a success rate higher than $95\%$. Furthermore, to be compatible with the setup in CortexBench authors, in particular, to keep the same camera locations, we filter out the tasks where the goal is not always visible in the camera FOV. This leads to a set of $15$ unknown tasks that are quite different from the tasks in the training set $T^{k}$ as they involve different objects (handle press, faucet, plate, door, window, etc.) and types of manipulation (sliding an object, lowering a press, opening a window, etc.). Each collected demonstration is a sequence of visual frames, proprioception inputs, and expert actions. The optimization of the task embedding is performed independently for each task (cf. §3.4). We use the AdamW optimizer and a learning rate of $1e{-1}$ during the task embedding search.
@@
Figure 6 presents the per-task performance in this setting. Despite the large variations between the new tasks in $T^{u}$ and the ones in the training set ($T^{k}$), the multi-task policy can adapt to many of them, without requiring any weight finetuning. Interestingly, the method performs particularly well on the Drawer Close task, which could be related to the presence of the time-inversed Drawer Open task in the training set. This provides some evidence that the method can exploit regularities between tasks, which seem to be captured by the task embedding space, making it possible to generalize to unseen variations. Figure 7 (a) shows qualitative examples of successful rollouts on the new tasks. The policy manipulates new objects (faucet, plate, window) and performs new moves (rotating the faucet or sliding the plate) not seen during training.
@@
Finally, Figure 7 (b) shows failure cases on unknown tasks. As seen on the first row, the policy avoids the wall, reaches the button, and starts pushing it, but fails to push it fully. This particular behavior was also observed on other rollouts, explaining the low success rate on this button-press-wall task while mastering a part of the required sub-skills. On the second row, the policy is able to move the cube but fails to bring it to the goal location (green dot). This gives some indication of the difficulty of the few-shot generalization case: exploiting regularities in the task space requires that tasks be performed more than just approximately, as often the success metric is sparse, and rollouts only count to the metric when they are executed fully and correctly.
Known task